AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Sat Feb 12 19:48:38 2022
Author
*
:
Subject
*
:
<p>> Initial grounding see final grounding from elog 187. ASIC settings for all - LEC slow comparator 0x64. Note on 11/02/22 1.8*L widths with same settings were aida1-4: 300,240,250,260</p> <p>ADC data items <10k</p> <p>Good events fluctuate 10k-20k</p> <p>1.8*L pulser widths aida1-4: 433,400,375,375(double peak)</p> <p>Attachments 1-4</p> <p>> New grounding implemented. 3mm grounding cable connect fee body to adaptor board. Each adaptor board is then connected directly to the cryring ground. Attachements 10-11</p> <p>ADC data items <10k</p> <p>Good events fluctuate 10k-20k</p> <p>1.8*L pulser widths aida1-4: 340,288,288,295 (reduced double peak)</p> <p>Noise in the system also seems to have decreaed in zoomed out 1.8*L plots (not attached). Aida1 still shows large noise peak channel 32000 on 1.8*L. 1.8*W no significant change, perhaps slightly reduced.</p> <p>Attachements 5-9</p> <p>Screencaps for following movement tests not taken as 1.8*L and 1.8*W spectra appeared the very similar as 5-9 except for slightly different widths. </p> <p>> Detectors moved in and out (pneumatic). Leakage current dropped by ~0.2 uA, torque now reads +2.8 previously was ~0. </p> <p>no change to ADC data items or good events. During movement aida3 saw momentary increase in good events to ~40k before returning to normal range. Rest of fees good events stats unaffected by movement.</p> <p>1.8*L pulser widths aida1-4: 413,313,295,306</p> <p>> servo motor (only) moved, torque reduced back to ~0</p> <p>no change in good events stats for any fees during servo movement. </p> <p>1.8*L pulser widths aida1-4: 425,335,316,288</p> <p>> Detector moved to IN position (pnuematic and servo). 19.3mm to 81.8mm on potentiometer. </p> <p>No change to good events. During movements aida3 saw momentary increase in good events to ~41k, all other fees unaffected. </p> <p>1.8*L pulser widths aida1-4: 302,221,305,293 </p> <p>Noise peak at channel ~32000 on aida 1 (1.8*L) significantly reduced. little noise on other fees </p> <p>>Detector moved to fully OUT position (pneumatic and servo). </p> <p>no change in good events, during movement aida3 saw momentary increase to ~35k</p> <p>1.8*L pulser widths aida1-4: 450,350,316,290</p> <p>Noise peak at chn 32000 on aida 1 has increased to previous size. little noise on other fees. </p> <p>> Bias increased to 120V. Leakage current stable however slightly higher ~0.1 uA</p> <p>no change in good events. </p> <p>1.8*L pulser widths aida1-4: 447,345,189,301</p> <p>> Bias increased to 140V, increase in leakage current ~0.2 uA, appears stable. Increase to 150V, fluctuate in leakage current of +/-0.5uA then huge increase to >20uA trip. </p> <p>> Fee power cycle. Bias 100V, leakage current 9.9 uA ~0.1-0.2 uA higher than previous but stable. </p> <p>no change to good events</p> <p>1.8*L pulser widths aida1-4: 468, 350,311, 296</p> <p>> Some decrease seen with new grounding, however main source of noise still apparent and any decrease seems insignificant compared to changes due to position movement. </p> <p>> Water pipes have been zipped tied so they no loger touch motors, no change in pulser widths or noise levels however. </p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Resubmit as new entry
Attachment 1:
2022-02-12_14-09-31aida1-2-5-1.8*W.png
Attachment 2:
2022-02-12_14-08-06aida3-4-1.8*W.png
Attachment 3:
2022-02-12_14-14-13aida1-2-1.8*L.png
Attachment 4:
2022-02-12_14-15-32aida3-4-1.8*L.png
Attachment 5:
2022-02-12_15-53-15aida1-2-1.8*W.png
Attachment 6:
2022-02-12_15-51-47aida3-4-1.8*W.png
Attachment 7:
2022-02-12_16-02-08aida1-2-1.8*L.png
Attachment 8:
2022-02-12_16-05-04aida3-4-1.8*L.png
Attachment 9:
2022-02-12_15-54-34aida1-4-rate.png
Original size: 1920x1080
Attachment 10:
20220212_152215.jpg
Original size: 2250x3000
Attachment 11:
20220212_171240.jpg
Original size: 2250x3000
Attachment 12:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown